Answer the following questions related to the given electrochemical cell under standard conditions.

W3+(aq) + 3e− ⇌ W(s)
E° = 0.100 V
Co2+(aq) + 2e− ⇌ Co(s)
E° = -0.280 V

1. Answer the following questions.

(a) The half cell containing W/W3+ is the  anode/ cathode

(b) Which one of the following statements isTRUE for the half cell containing W and W3+.
W3+ will be oxidized to form W.
W3+ will be reduced to form W.
W will be reduced to form W3+.
W will be oxidized to form W3+.

(c) The half cell containing Co2+/Co is the  anode/ cathode

(d) Which one of the following statements isTRUE for the half cell containing Co and Co2+.
Co will be reduced to form Co2+.
Co will be oxidized to form Co2+.
Co2+ will be reduced to form Co.
Co2+ will be oxidized to form Co.


Tries 0/3

2. What is E°cell (in V)? Report your answer to three decimal places in standard notation (i.e., 0.123 V).


Tries 0/3

3. Answer the following questions.

(a) Which one of the following is the balanced chemical equation for the spontaneous process occurring in the electrochemical cell under standard conditions?
3W3+(aq) + 2Co(s) → 3W(s) + 2Co2+(aq).
2W(s) + 3Co2+(aq) → 2W3+(aq) + 3Co(s).
3W(s) + 2Co2+(aq) → 3W3+(aq) + 2Co(s).
2W3+(aq) + 3Co(s) → 2W(s) + 3Co2+(aq).

(b) How many electrons are being transferred in this balanced chemical reaction?

(c) What is ΔG° (in kJ/mol)? Report answer to three significant figures in scientific notation (i.e., 1.23e2 kJ/mol)


Tries 0/3

4. If the reaction was allowed to proceed, the mass of the W electrode will (increase /decrease ) and the mass of the Co electrode will (increase/ decrease ) . The [W3+] in solution will ( increase /decrease) and the [Co2+] in solution will (increase/ decrease)


Tries 0/3

5. The salt bridge contains K+ and Cl. As the cell is in operation the K+ ions will tend to migrate toward the (anode /cathode) and the Cl ions will tend to migrate toward the ( anode/ cathode) . This occurs to counter balance any loss of ( W^3+ /Co^2+ ) ions as they precipitate onto the (W /Co) electrode


Tries 0/3
Answer #1

W3+(aq) +3e- = W(s), Here Reduction Occurs as W3+ is oxidized (loses 3 electrons) to form W(s).

The half-reaction on the cathode where reduction occurs. also there is +ve potential value E° = 0.100 V

Co(s) = Co2+(aq) + 2e , Here Oxidation Occurs as Co(s) is reduced (gain 2 electrons) to form Co2+(aq)

The half-reaction on the anode where Oxidation occurs. there is a negetive potential value E° = -0.280 V

Balance eqution,

W3+(aq) + Co(s) = W(s)+  Co2+(aq)

3W3+(aq) + 2Co(s) → 3W(s) + 2Co2+(aq).

Ecell=Ecathode-Eanode = 0.100 V-(-0.280 V)=0.380V